# KG-IRAG: A Knowledge Graph-Based Iterative Retrieval-Augmented Generation Framework for Temporal Reasoning

### v2

@@ 3.2 Knowledge Graph Construction
[…]
9:     $\text{CombinePlan}(\cdot)=\text{CPPlan}$
[…]
11:end Declare
12:Algorithm Begin
13:$k\leftarrow\text{CP}(time_{i},loc_{i},RP)$
14:$judge\_result\leftarrow\text{LLM2.JS}(k)$
15:while $judge\_result$ is not sufficient do
16:     $event_{abnormal}\leftarrow\text{FAE}(time_{i},loc_{i})$
17:     $time_{i^{\prime}},location_{i^{\prime}}$triangleright Generate next time period or locations based on abnormal events in current time slots or locations
18:     if $time_{i^{\prime}},location_{i^{\prime}}$ exists then
19:         $\text{CTL}(time_{i},loc_{i},time_{i^{\prime}},location_{i^{\prime}})$
20:         $k\leftarrow\text{CP}(time_{i},loc_{i},RP)$
21:         $judge\_result\leftarrow\text{LS}(k)$
22:     else
23:         return "no answer" $\triangleright$ If no more time and location can be retrieved
24:     end if
25:end while
26:if $judge\_result$ is sufficient then
27:     $final\_plan\mathrel{\leftarrow}\text{CPPlan}(time_{i},loc_{i})$
28:     $final\_triplets\mathrel{\leftarrow}\text{KGret}(final\_plan)$
29:     Generate $RAG\_Data\_Prompt$
30:     Combine $RP$ and $RAG\_Data\_Prompt$
31:     return $answer$
32:end if
33:Algorithm End
3.3 Triplet Retrieval
Once the starting time (start time 0) and location (location 0) have been identified by LLM1, KG exploration begins. The retrieval process involves searching the knowledge graph for relevant triplets that match the query’s initial time and location. This process is guided by LLM1’s reasoning prompt, which provides a high-level explanation of the information needed to solve the query. In each iteration, the system retrieves a set of triplets (triplets1) that are relevant to the current query conditions. The retrieval is conducted through a series of KG searches that target both time and spatial relationships, ensuring that all possible relevant data is retrieved for the given query. The retrieved triplets are then passed to LLM2, which assesses whether the current data, combined with the reasoning prompt, is sufficient to answer the query. If the data is insufficient, the system moves to the next time slot or location, retrieving a new set of triplets (triplets2, triplets3, etc.) until LLM2 determines that the question can be resolved.
3.4 Iterative Reasoning
The iterative reasoning process is the core mechanism of KG-IRAG for handling complex temporal queries. After each retrieval of triplets, the system evaluates whether current set of triplets, combined with the reasoning prompt generated by LLM1, is sufficient to answer the query. LLM2 is responsible for this evaluation, determining whether the current evidence from the KG provides a valid solution or if further exploration is needed. If LLM2 deems that the current triplets are insufficient to resolve the query, the system yields to a different time or location, refining the search parameters. For example, if the goal is to avoid rain during a trip, and the system detects rainfall at 9:00 AM, it may retrieve data from earlier or later time slots to identify an optimal departure time. This iterative process continues until the system finds a time and location that satisfy the query. Algorithm 2 outlines the step-by-step process of iterative reasoning for temporal queries, where LLM1 directs the search and LLM2 evaluates the retrieved data to determine retrieval should be continued.
3.5 Answer Generation
Once LLM2 confirms that a sufficient set of triplets has been retrieved to resolve the query, the final step is to generate the answer. The triplets from the various iterations of the KG search are combined into a coherent RAG Data Prompt, and the sufficient triplet evidences are then processed together with the former reasoning prompt, together to get the answer generated. If data is insufficient to yield the answer or is out of boundary during KG lookup, ‘no answer’ is returned as the result.
3.6 Case Study
To further explain the details of ‘iteration‘ in KG-IRAG, a case study was conducted. Presented in section A, it simulate an initial travel plan to Sydney Opera House, including defining whether the initial plan is valid, as well as how to define an optimal time to adjust the trip.
4 Dataset properties
4.1 Data properties
To further test a RAG model’s ability to explore and do logical reasoning on temporal data, 3 datasets, weatherQA-Irish, weatherQA-Sydney and trafficQA-TFNSW are prepared for experiments. From the perspective of RAG, the construction of these datasets is tailored to allow the model to perform both entity-based retrieval and time-dependent reasoning. beginning with straightforward entity identification and advancing to dynamic temporal reasoning over multiple steps. Sydney weather data is comprised of half-hourly records, while Irish weather and TFNSW data are collected on an hourly basis. The high-resolution temporal granularity of these datasets allows for the construction of knowledge graphs that not only record spatial relationships but also capture detailed temporal data. What’s more, in the TFNSW dataset, there are also traffic directions, offering more details of spatial information. In each dataset, attributes such as date, location, and event status (e.g., rainfall or traffic volume) are structured as KG entities and relations.
4.2 Question Designs
The design of the questions in the QA datasets is structured to progressively increase the complexity of the retrieval tasks. The first question (Q1) is a fundamental entity recognition and retrieval task, while the second (Q2) and third (Q3) questions introduce logical reasoning by incorporating time-dependent queries. The goal is to examine the system’s ability to not only retrieve relevant data but also engage in iterative reasoning over time.
4.2.1 Question 1: Abnormal Event Detection
Q1 is designed as a straightforward retrieval task in which the model identifies whether an "abnormal" event, such as rainfall or traffic congestion, occurred during a specific time slot. This task relies primarily on entity recognition and retrieval of static information from the knowledge graph. As a result, it reflects a typical RAG problem, where the system extracts relevant data based on predefined entities and times.
4.2.2 Questions 2 and 3: Leave Early or Late
In contrast to Q1, the design of Q2 and Q3 introduces temporal reasoning components that significantly increases the complexity of the retrieval process. Q2 asks the model to calculate the latest time at which one can leave early to avoid an abnormal event, while Q3 focuses on determining the earliest time to leave late. Both questions require the model to infer optimal departure times by reasoning over a temporal range, making use of the temporal relationships embedded in KGs.
From the RAG perspective, Q2 and Q3 can be seen as decompositions of multiple Q1-type sub-questions. Instead of a single, static retrieval, the model must query the knowledge graph iteratively, retrieving data from different time slots and integrating that information to form a cohesive answer. This process illustrates the key idea of KG-IRAG: transforming a larger temporal question into multiple fixed-time subproblems, each solvable via entity and temporal information extraction.
@@ 4.2.3 Dynamic Problem Decomposition
The design of Q2 and Q3 exemplifies the concept of dynamic problem decomposition in retrieval-augmented generation. By requiring the model to engage in time-based reasoning, these questions push beyond standard entity recognition and challenge the system to handle a broader scope of temporal logic. LLMs need to solve the query through a combination of logical steps.
@@ 5 Experiments
5.1 Datasets
Three datasets are used to evaluate the performance of KG-IRAG: Irish weather data, Sydney weather data, and Traffic Volume of Transport for New South Wales (TFNSW) data. While the two weather datasets are used to identifying abnormal events like rainfall, the TFNSW dataset includes extensive numerical data on traffic volume, adding complexity to the questions by requiring numerical comparisons. All three datasets are transferred into KGs that capture location relationships and temporal records. While the details od datasets are presented in B,Table 1 presents attributes used to form the KGs, with time treated as an entity to enhance retrieval capabilities. Table 2 represents the size of each dataset and the corresponding KGs generated.
To form the QA datasets, for each year, 200 time slots were chosen randomly, and the questions are generated based on the following criteria: Question 1 asks about in the chosen time slot, whether there is certain ’abnormal’ activity(i.e. rainfall, traffic jam), answer with ’True’ or ’False’. If the answer is yes in question 1, question 2 and 3 care about what is the latest time to head off early and the earliest time to head off late under the same length of the trip. (the maximum early or late time is set to be 12 hours for weather datasets and 9 hours for the TFNSW dataset). Table 3 shows the details of question proportion in the three datasets.
@@ 5.2 Evaluation
Experiments are conducted to evaluate the datasets. For each question, a minimal subset of data, referred to as "standard data," is selected according to the criteria outlined in equation (1) and (2).
The standard data represents the minimal amount of information necessary to answer the query, ensuring that only the relevant data is used (e.g., if data from 7:00 to 12:00 is required to answer a question related to the status of 7:00 to 10:00, then no additional data outside 7 to 12 is included). Comparisons are made by feeding the standard data into LLMs in raw data format (data frame), context-enhanced data, and KG triplet representations.
@@
Standard evaluation metrics are employed to assess QA systems: Exact Match (EM), F1 Score, and Hit Rate (HR). In addition, hallucinations are considered to evaluate the accuracy of KG-IRAG compared with other methods. These metrics collectively measure the model’s capability to provide accurate answers while avoiding the retrieval of both excessive and insufficient knowledge. Exact Match (EM) calculates the percentage of results that exactly match the ground truth answer, defined as $EM(\%)=\frac{\Sigma_{i=0}^{N}answer_{i}=truth_{i}}{N}$ F1 Score is calculated based on precision and recall. Precision measures the correctness of the retrieved information, while recall assesses how much of the target data has been retrieved, shown as $Precision=\frac{I_{retrieved}\cap SD}{I_{retrieved}}$ , $Recall=\frac{I_{retrieved}\cap SD}{SD}$ and $F1=\frac{2*Precision*Recall}{Precision+Recall}$
@@
Hallucination refers to content generated by LLMs that is not present in the retrieved ground truth Ji et al. (2023); Li et al. (2024); Perković et al. (2024). It is a crucial metric for assessing the accuracy and reliability of QA systems, calculated as shown in (4). Specifically, hallucination is defined as follows during the evaluation: For Question 1, hallucination occurs if the answer is incorrect, as this indicates that the LLM failed to correctly identify the abnormal event despite the relevant data being provided. For Questions 2 and 3, hallucination is detected if the LLM: (a) generates an answer indicating an abnormal event at an incorrect time; (b) produces an answer that does not appear in the provided data; or (c) in the case of the KG-IRAG system, LLM2 fails to decide when to stop the exploration process.
@@
While hallucination can be directly calculated for Question 1, whereas for Questions 2 and 3, 50 questions are randomly selected from each dataset and manual reviews are conducted to assess the hallucinations in the answers from LLMs. To benchmark the performance of the proposed model (KG-IRAG), several LLMs are utilized Touvron et al. (2023); Achiam et al. (2023): 1) Llama-3-8B-Instruct, 2)GPT-3.5-turbo-0125 3) GPT-4o-mini-2024-07-18, 4)GPT-4o-2024-08-06. KG-IRAG is compared against two RAG methods: The first one is a standard one without exploration, by asking LLMs to decide the data needed to solve the problem, retrieving the data then feed into LLMs. The second one is KG-RAG, which contains Chain of Exploration for KG Retrieval Sanmartin (2024).
@@ 6.1 Experiment Results
6.1.1 Testing LLMs using Facts
The first set of experiments focuses on testing LLMs directly on the datasets. Standard data are calculated and collected according to equation  (1) and  (2). This data is converted into three formats: raw data (in table format), text data (by converting the data into text descriptions in various forms), and triplet data (extracted from knowledge graphs). The experiments are designed to assess whether LLMs can effectively use these different data formats, and which format is best recognized by LLMs. For each question, the different formats of the standard data are combined with the query into a single prompt. The goal is to determine whether LLMs can solve problems involving temporal and logical reasoning without iterative reasoning but with the correct data. The results are shown in table 4.
6.1.2 Testing different RAG systems
The second stage compares the performance of three different RAG systems: 1)Standard Graph-RAG system, where LLMs are prompted to determine "what data is needed to solve the problem," retrieve the necessary data based on responses and then provide the final answer; 2)KG-RAG Sanmartin (2024), which uses a Chain of Exploration to retrieve data step-by-step, with the exploration plan set to three steps, considering the temporal span in the questions; 3)KG-IRAG: our proposed system. For testing KG-IRAG, the same LLM1 and LLM2 models (as illustrated in Figure 2) are used for both data retrieval and iterative reasoning. Detail results of exact match, F1 score, hit rate and hallucination for different models applying RAG systems on the 3 datasets are shown in table 5,  6, and  7, while the plot of exact match is in figure 4.
@@ 6.2 Result analysis
6.2.1 Different Format of Data Input
@@ 6.2.2 RAG Comparation: Question 1
For Q1, all three systems (Standard Graph-RAG, KG-RAG, KG-IRAG) perform similarly. This is because LLMs do not require step-by-step reasoning to determine which data to retrieve from the database—they can reliably identify entity names and times in the queries and retrieve correct data through one QA. Consequently, the results for Question 1 are comparable to simply feeding correct data to the LLMs and asking them to detect abnormal activities. Therefore, no significant differences were observed between the three systems in this task.
@@ 6.2.4 Hallucinations
1. ‘Late Stop‘ Phenomenon: During evaluation, KG-IRAG occasionally tends to stop late and continues retrieving data unnecessarily. Specifically, in manny cases LLM2 in KG-IRAG decided to retrieve one or two more round of data, leading to an overload of information. The ‘hallucination‘ occured more easily in questions which contain too many numbers, like the TrafficQA-TFNSW.
2. In other baselines, hallucination often lead to ‘LLMs generate wrong answer using insufficient data‘, which is considered more harmful since it directly leads to wrong conclusion.
3. Although KG-IRAG indeed tend to retrieve some extra rounds of data, that hallucination won’t normally lead to a wrong answer, since sufficient data is already in the retrieval plan.
@@ 7 Conclusions
In this paper, a new RAG framework is proposed, i.e., Knowledge Graph-Based Iterative Retrieval-Augmented Generation(KG-IRAG), which enhances the integration of KG with LLMs through iterative reasoning and retrieval. KG-IRAG allows LLMs to make more informed decisions when answering complex, temporally dependent queries by progressively retrieving relevant data from KGs. Unlike traditional RAG methods, KG-IRAG flexibly and effectively employs a step-by-step retrieval mechanism that guides LLMs in determining when to stop exploration, significantly improving response accuracy.
To evaluate the framework’s effectiveness, three new datasets, weatherQA-IrishweatherQA-Sydney, and trafficQA-TFNSW, have been introduced. These datasets, which involve real-world scenarios such as weather and traffic conditions, are designed to test the ability of LLMs to handle time-sensitive and event-based queries requiring both temporal reasoning and logical inference. Experimental results demonstrate that KG-IRAG excels in complex reasoning tasks by generating more accurate and effecient responses.
@@ 8 Limitations
Although KG-IRAG presents notable improvements in LLMs’ ability to tackle knowledge-intensive tasks, there are still several limitations, including:
(1) There is a need for further enhancement in the reasoning mechanism to ensure more accurate and efficient data management.
(2) The influence of imperfect retrieval(retrieve extra data) in some cases remains a challenge, and future improvements should focus on mitigating its impact on performance.
(3) There would be potential need for balancing the use of internal and external knowledge, such as Wang et al. (2024)
9 Acknowledgements
This research is partially supported by the Technology Innovation Institute, Abu Dhabi, UAE. Additionally, computations were performed using the Wolfpack computational cluster, supported by the School of Computer Science and Engineering at UNSW Sydney.
Appendix A Case study

### v3

@@ 4.2.2 Questions 2 and 3: Leave Early or Late
10:         $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{(q_{2},t_{\text{early}}),(q_{3},t_{\text{late}})\}$
11:     end if
12:end for
@@ 4.2.3 Dynamic Problem Decomposition
From the RAG perspective, Q2 and Q3 can be seen as decompositions of multiple Q1-type sub-questions. Instead of a single, static retrieval, the model must query the knowledge graph iteratively, retrieving data from different time slots and integrating that information to form a cohesive answer. This process illustrates the key idea of KG-IRAG: transforming a larger temporal question into multiple fixed-time subproblems, each solvable via entity and temporal information extraction. By requiring the model to engage in time-based reasoning, these questions push beyond standard entity recognition and challenge the system to handle a broader scope of temporal logic. LLMs need to solve the query through a combination of logical steps.
@@ 5.1 Evaluation
KG-IRAG is evaluated on 3 datasets mentioned above. For each question, a minimal subset of data, referred to as "standard data," is selected according to the criteria outlined in equation (15) and (16). The standard data represents the minimal amount of information necessary to answer the query, ensuring that only the relevant data is used (e.g., if data from 7:00 to 12:00 is required to answer a question related to the status of 7:00 to 10:00, then no additional data outside 7 to 12 is included). Comparisons are made by feeding the standard data into LLMs in raw data format (data frame), context-enhanced data, and KG triplet representations.
@@
Standard evaluation metrics are employed to assess QA systems: Exact Match (EM) calculates the percentage of results that exactly match the ground truth answer, defined as $EM(\%)=\frac{\Sigma_{i=0}^{N}answer_{i}=truth_{i}}{N}$ F1 Score is calculated based on precision and recall. Precision measures the correctness of the retrieved information, while recall assesses how much of the target data has been retrieved, shown as $Precision=\frac{I_{retrieved}\cap SD}{I_{retrieved}}$ , $Recall=\frac{I_{retrieved}\cap SD}{SD}$ and $F1=\frac{2*Precision*Recall}{Precision+Recall}$
@@
Hallucination refers to content generated by LLMs that is not present in the retrieved ground truth Ji et al. (2023); Li et al. (2024a); Perković et al. (2024), calculated as shown in (18). Specifically, hallucination is defined as follows during the evaluation: For Question 1, hallucination occurs if the answer is incorrect, as this indicates that the LLM failed to correctly identify the abnormal event despite the relevant data being provided. For Questions 2 and 3, hallucination is detected if the LLM: (a) generates an answer indicating an abnormal event at an incorrect time; (b) produces an answer that does not appear in the provided data; or (c) in the case of the KG-IRAG system, LLM2 fails to decide when to stop the exploration process.
@@
While hallucination can be directly calculated for Question 1, whereas for Questions 2 and 3, 50 questions are randomly selected from each dataset and manual reviews are conducted to assess the hallucinations in the answers from LLMs. To benchmark the performance of the proposed model (KG-IRAG), several LLMs are utilized Touvron et al. (2023); Achiam et al. (2023): 1) Llama-3-8B-Instruct; 2)GPT-3.5-turbo-0125; 3) GPT-4o-mini-2024-07-18; 4)GPT-4o-2024-08-06. KG-IRAG is compared against two RAG methods: The first one is a standard one without exploration, by asking LLMs to decide the data needed to solve the problem, retrieving the data then feed into LLMs. The second one is KG-RAG, which contains Chain of Exploration for KG Retrieval Sanmartin (2024).
5.2 Experiment Settings
5.2.1 Dataset Verifications
Standard data are calculated and collected according to equation  (15) and  (16), then converted into three formats: $d_{raw}$ (in table format), $d_{text}$ (by converting $d_{raw}$ into text descriptions in various forms), and $d_{triplet}$ (extracted from knowledge graphs). The experiments are designed to assess whether LLMs can effectively answer questions, and which format is best recognized by LLMs. For each question $\mathcal{Q}$, $\{d_{raw},d_{text},d_{triplet}\}$ are inserted into template to form prompt. The goal is to determine whether LLMs can do temporal reasoning on our datasets without iterative reasoning but with the correct data. The results are shown in table 5.
5.2.2 KG-IRAG Verifications
The second stage verifies KG-IRAG with baselines: 1)Single-pass Graph-RAG system, where LLMs are prompted to determine needed data once and retrieving the necessary data based on responses and then provide the final answer; 2)KG-RAG Sanmartin (2024), which uses a Chain of Exploration to retrieve data step-by-step, with the exploration plan set to three steps, considering the temporal span in the questions; 3)KG-IRAG: our proposed system. For testing KG-IRAG, the same LLM1 and LLM2 models (as illustrated in Figure 2) are used for both data retrieval and iterative reasoning. Detail results of exact match, F1 score, hit rate and hallucination for different models applying RAG systems on the 3 datasets are shown in table 6,  7, and  8, while the plot of exact match is in figure 3.
@@ 6.1 Result analysis
6.1.1 Testing LLMs on our datasets
@@ 6.1.2 RAG Comparation: Question 1
For Q1, all three systems (Standard Graph-RAG, KG-RAG, KG-IRAG) perform similarly. This is because LLMs do not require step-by-step reasoning to determine which data to retrieve from the database—they can reliably identify entity names and times in the queries and retrieve correct data through one QA. Consequently, the results for Question 1 are comparable to simply feeding correct data to the LLMs and asking them to detect abnormal activities. Therefore, no significant differences were observed between the three systems.
@@ 6.1.4 Hallucinations
1. KG-IRAG: ‘Late Stop‘ Phenomenon: During evaluation, KG-IRAG occasionally tends to stop late and continues retrieving data one or two nodes deeper in the path. The ‘hallucination‘ occured more easily in questions which contain too many numbers, like the TrafficQA-TFNSW.
2. Baselines: Insufficient or Fabrication: In other baselines, hallucination often lead to ‘LLMs generate wrong answer using insufficient data or even fake data‘, causing lower accuracy.
3. Comparison: As proven in Appendix B.1, KG-IRAG ensures higher accuracy with least chance of getting useless data at the beginning. The hallucination, although makes a larger dependency depth $d_{kgIRAG}>\hat{d}$, proves KG-IRAG is retrieving data through correct and single evidence path, while other baselines fall into chaos earlier in depth $d_{b}<\hat{d}$ and end in depth $d_{e}>\hat{d}$, with probability of missing correct evidence path decaying exponentially as $\delta^{-(d_{b}-d_{e})}=e^{-\lambda(d_{b}-d_{e})}$, shown in Equation 19, causing both waste in tokens and higher latency to correct retrieval.
@@ 7 Conclusions
In this paper, a new RAG framework is proposed, i.e., Knowledge Graph-Based Iterative Retrieval-Augmented Generation(KG-IRAG), which enhances the integration of KG with LLMs through iterative reasoning and retrieval. Unlike traditional RAG methods, KG-IRAG flexibly and effectively employs a step-by-step retrieval mechanism that guides LLMs in determining when to stop exploration, significantly improving response accuracy. To evaluate the framework’s effectiveness, three new datasets, weatherQA-IrishweatherQA-Sydney, and trafficQA-TFNSW, have been introduced. These datasets involve real-world scenarios, designed to test the ability of LLMs to handle time-sensitive and event-based queries requiring both temporal reasoning and logical inference. Experimental results demonstrate that KG-IRAG excels in complex reasoning tasks by generating more accurate and efficient responses.
@@ 8 Limitations
Although KG-IRAG presents notable improvements in LLMs’ ability to tackle knowledge-intensive tasks, there are still several limitations, including: (1) Further enhancement in the reasoning mechanism to ensure more accurate and efficient data management; (2) Potential need for balancing the use of internal and external knowledge, such as Wang et al. (2024)
Appendix A Notations
Important notations are listed in Table 1.
Table 1: Notations Tables in KG-IRAG
Notation Definition $\mathcal{F}$ Abbreviation of the KG-IRAG framework. $\mathcal{A}$ Temporal Anchoring Module. $\mathcal{C}$ Iterative Reasoning Controller $\Psi$ Retrieval Operator $D$ Knowledge Graph / Subgraph, where $D^{i}$ is result after $i$-th retrieval. $S^{t}$ Reasoning State at step(time) $t$ $\tau^{(t)},\mathcal{T}$ Temporal anchor point, current time point on KG $\mathcal{G}$ $\ell^{(t)}$ Location anchor point, current location point on KG $\mathcal{G}$ $\mathcal{R}^{(t)}$ Retrieved triplet at time $t$ $\mathcal{P}^{(t)}$ Reasoning prompt leading retrieval $\mathcal{M}_{1},\mathcal{M}_{2}$ LLM1 and LLM2 actions $\mathbb{I}_{\text{stop}}$ Termination condition deciding whether to stop whole process $K^{(t)}$ Number of hop for each iteration $\mathcal{Q}$ Questions, combining $\mathcal{Q}_{base}$ and $\mathcal{Q}_{reason}$ $SD$ Minimal subset of data needed to answer each question
Appendix B Theoretical Analysis of Iterative Retrieval
B.1 Algorithmic Necessity of Iteration
Let $\mathcal{P}$ be the class of temporal reasoning problems requiring multi-hop dependencies. By complexity analysis, cases exist when single-pass RAG cannot solve $\mathcal{P}$ efficiently:
Definition B.1 (Temporal Dependency Depth).
For problem $p\in\mathcal{P}$, define dependency depth $d(p)=\hat{d}$ as the minimum number of temporal hops required to connect anchor time $\tau_{0}$ to solution window $\tau^{*}$.
Lemma B.2 (Single-pass Success Probability).
Single-pass RAG with retrieval radius $k$ succeeds only if $k\geq\hat{d}$. For $p$ with unbounded $\hat{d}$, if single-pass retrieval returns all nodes within radius $k<d$,the probability that the evidence path is fully covered decays exponentially:
$P_{\text{success}}(k)\leq\delta^{-(\hat{d}-k)}=e^{-\lambda(\hat{d}-k)},\qquad\lambda=\ln\delta$ (19)
Proof.
Let retrieval cost $C(k)=|N_{k}(v_{0})|$. At depth k the subgraph tree has $\delta^{k}$ nodes. For KG with average degree $k$, $C(k)=O(\delta^{k})$. The evidence path’s terminal node lies uniformly at depth $\hat{d}$ For one of the $\delta^{\hat{d}}$ at depth $\hat{d}$, the retrieval misses it unless its ancestor at depth $k$ is retrieved. As the ancestor is unique and appear among $\delta^{k}$ nodes, the chance of hit comes with probability $\delta^{k}/\delta^{d}=\delta^{-(d-k)}$
Cost comparison.
Retrieval cost grows as $C(k)=|\mathcal{N}_{k}|=\Theta(\delta^{k})$, so using $k=\hat{d}$ is often infeasible for large $\hat{d}$. Iterative retrieval instead performs $T$ queries of radius $k_{0}\ll\hat{d}$, incurring total cost $T\cdot\Theta(\delta^{k_{0}})$ with $T=\lceil\hat{d}/k_{0}\rceil$, which is asymptotically cheaper when $\hat{d}\to\infty$.
B.2 Convergence Analysis
Theorem B.3 (Accuracy Preservation).
Let $\mathcal{S}^{*}$ be the set of relevant facts and $\mathcal{S}_{t}$ the facts retrieved at iteration $t$. Denote precision $p_{t}=\frac{|\mathcal{S}^{*}\cap\mathcal{S}_{t}|}{|\mathcal{S}_{t}|}$ and single-pass precision $p_{\text{single}}=\frac{|\mathcal{S}^{*}|}{|\mathcal{S}_{\text{single}}|}$. If each iterative query is restricted to radius $k_{0}<\hat{d}$, then $p_{t}\;\geq\;\delta^{-(k_{0}-\alpha)}$, $p_{\text{single}}=\delta^{-(\hat{d}-\alpha)}$, and $\alpha:=\log_{\delta}|\mathcal{S}^{*}|$. So $p_{t}>p_{\text{single}}$ whenever $k_{0}<\hat{d}$. Consequently,
$\prod_{t=1}^{T}p_{t}\geq p_{\text{single}}^{T}$ (20)
Proof.
Because each iterative step retrieves at most $\delta^{k_{0}}$ nodes, $|\mathcal{S}_{t}|\leq\delta^{k_{0}}$. Substituting into the definition of $p_{t}$ and comparing with $p_{\text{single}}$ yields the inequalities above. ∎
Appendix C Case study